# A Regression-Based Approach to the CO2 Airborne Fraction: Enhancing Statistical Precision and Tackling Zero Emissions

### v3

@@ Abstract
Previous studies have analyzed the AF as the ratio of yearly changes in atmospheric CO2 concentrations ($G_{t}$, numerator) and anthropogenic CO2 emissions ($E_{t}$, denominator) (e.g. Canadell et al., 2007; Raupach et al., 2008; Le Quéré et al., 2009; Knorr, 2009; Ballantyne et al., 2012; Raupach et al., 2014; Ballantyne et al., 2015; Keenan et al., 2016; Bennedsen et al., 2019; van Marle et al., 2022; Pressburger et al., 2023). Following this literature, the conventional estimator of the AF has been defined as the sample mean of the yearly ratio $G_{t}/E_{t}$. In this paper, we show that this estimator suffers from a number of statistical deficiencies due to its definition as the ratio of two stochastic processes. One problem is that the denominator in the ratio, $E_{t}$, may have a positive probability density at zero, implying that the estimator does not possess any moments. This lack of moments implies, for instance, that the mean and variance of the estimator do not exist. Although this issue may be of limited concern during the period 1959–2022, it becomes important in future scenarios where CO2 emissions decrease, such as those consistent with “net-zero” CO2 emissions, a committed goal of the international community (Rockström et al., 2017; Riahi et al., 2022). For example, issues with the conventional analysis of the AF when emissions decrease are recently encountered in Pressburger et al. (2023), where the future AF implied by output from a climate model is studied. Another problem is that the time series of changes in atmospheric concentrations $G_{t}$ and emissions $E_{t}$ exhibit a trend, which implies that a central limit theorem does not hold for the estimator. Therefore, its limiting distribution may be non-Gaussian, unless a separate assumption of Gaussianity is imposed. This implies that confidence intervals and $p$-values for test statistics based on the Gaussian distribution may not be valid for the conventional ratio-based estimator.
@@
We show that the time series of changes in atmospheric CO2 concentrations, $G_{t}$, cointegrates with the time series of anthropogenic CO2 emissions, $E_{t}$, over the period 1959–2022. Cointegration means that, even though $G_{t}$ and $E_{t}$ are non-stationary time series individually, the linear regression equation $G_{t}=\alpha E_{t}+u_{t}$, with constant parameter $\alpha$, has a stationary error $u_{t}$. In other words, the non-stationary processes $G_{t}$ and $E_{t}$ are cointegrated when a linear combination exists that yields a stationary process. The parameter $\alpha$ denotes the fraction of emissions that is added to the atmosphere in year $t$, i.e., in the linear regression, $\alpha$ represents the AF. We show the constancy of $\alpha$ statistically on the historical sample. In future scenarios where emissions decrease substantially, $\alpha$ will be time-varying, however.
@@
Figures 1a) and 1b) show yearly changes in atmospheric CO2 ($G_{t}$) and yearly CO2 emissions from anthropogenic sources ($E_{t}$), respectively. The black line in Figure 1c) shows the ratio of these two variables, $G_{t}/E_{t}$. Data are obtained from the Global Carbon Project and cover the period 19591959195919592022202220222022; in the empirical section below we provide further details on the data. The most conspicuous feature of the two data series $G_{t}$ and $E_{t}$ is that both time series exhibit upwards trends. Trending behavior is indicative of time series being non-stationary. A simple least-squares statistical analysis of the bivariate system $(G_{t},E_{t})$, where the non-stationarity of a time series is not accounted for, yields invalid inference and should be avoided (Granger and Newbold, 1974). However, the notion of cointegration (see, for example, chapter 19 in Hamilton, 1994, for a textbook treatment) allows us to keep working with the trending time series $G_{t}$ and $E_{t}$ while still obtaining valid statistical inference. Cointegration methods have been applied in earlier climate studies (e.g. Kaufmann and Stern, 2002; Schmith et al., 2012). Informally, two time series are cointegrated if they share a common trend. Formally, the time series $G_{t}$ and $E_{t}$ are said to be cointegrated when both $G_{t}$ and $E_{t}$ are non-stationary, and the error term $u_{t}$ in the regression equation $G_{t}=\alpha E_{t}+u_{t}$ is stationary. We adopt the Dickey-Fuller test (Dickey and Fuller, 1979) to determine whether a time series is non-stationary. The test statistic is for the null hypothesis of a unit root, that is, of having a unit value for the autoregressive dependence of $G_{t}$ on its lagged value $G_{t-1}$ (and $E_{t}$ on $E_{t-1}$). The results from this test strongly suggest that both time series are non-stationary (Supplementary Information S2.1). The null hypothesis of no cointegration (that is, $u_{t}$ is non-stationary) can be tested formally using the Engle-Granger test (Engle and Granger, 1987), which is a Dickey-Fuller test on the residuals in the regression $G_{t}=\alpha E_{t}+u_{t}$, adjusted for the fact that these residuals are not observed but must be estimated. The null hypothesis of a unit root in $u_{t}$ is firmly rejected (Supplementary Information S2.1), and we can therefore conclude that the two time series $G_{t}$ and $E_{t}$ are cointegrated.
@@
for $t=1,2,\ldots,T$, where the disturbance $u^{(2)}_{t}$ is also a zero-mean error process. The relationship between the disturbances in equations (1) and (2) is given by $u^{(1)}_{t}=u^{(2)}_{t}\,/\,E_{t}$. Cointegration of $G_{t}$ and $E_{t}$ implies that $u_{t}^{(2)}$ is a stationary process. Then, the parameter $\alpha$ can be estimated directly using a simple least-squares calculation, yielding the regression-based estimator as given by
@@
Table 1: Least-squares regression output of the linear models (1)–(4) for the Global Carbon Project data on the full sample 1959–2022 (left panel) and the subsample 1992–2022 (right panel). Estimates of the AF $\alpha$ are denoted by“$\hat{\alpha}$.” The table reports standard errors “SE($\hat{\alpha}$)” and standard errors relative to SE($\hat{\alpha}_{1}$) from model (1). Relative SE less than one indicates lower estimation uncertainty than the ratio-based estimator. The $95\%$ confidence interval for the AF $\alpha$, based on the Gaussian distribution, is denoted by “$CI_{95\%}(\alpha)$”, the estimated standard deviation of the error $u_{t}$ by “$\widehat{SD}(u_{t})$”, and the coefficient of determination by “$R^{2}$”.
@@
By introducing covariates in models (1) and (2), we can reduce the variance of the error processes $u_{t}$ and thus achieve more precise estimates of the AF $\alpha$. For example, it is common practice in the literature to control for the effects of El Niño ($ENSO_{t}$) and volcanic activity ($VAI_{t}$), (e.g. Raupach et al., 2008; van Marle et al., 2022). [^3] ENSO data is the Niño 3 SST Index from the National Oceanic and Atmospheric Administration (NOAA), available at https://psl.noaa.gov/gcos_wgsp/Timeseries/Data/nino3.long.anom.data, last accessed June 17, 2024, and data from volcanic activity is from Ammann et al. (2003). The ENSO data contain a slight trend, which was removed by de-trending before using the data in the analysis of this paper, to avoid having the trend influence the estimates of the AF. The ENSO and VAI data series are shown in Figure S2 in the Supplementary Information. Hence, we consider the models
@@
for $t=1,2,\ldots,T$, where $\tilde{\gamma}_{i}$ and $\gamma_{i}$, for $i=1,2$, are regression coefficients, and the model errors $u^{(j)}_{t}$ follow a zero-mean error process, for $j=3,4$. For both models, the coefficients can be estimated using least-squares regression. Let $\hat{\alpha}_{3}$ and $\hat{\alpha}_{4}$ denote the least-squares estimators of $\alpha$ from models (3) and (4), respectively. The estimation results relevant for the AF $\alpha$ are presented in columns labeled as (3) and (4) of Table 1, while the estimation results for the ENSO and VAI coefficients $\gamma_{i}$ and $\tilde{\gamma}_{i}$, for $i=1,2$, are reported in Table S2 of the Supplementary Information. The regression fits of models (3) and (4) are presented as in Figure 1c) and 1e) and their associated estimated residuals $\hat{u}_{t}$ in Figure 1d) and 1f), all as red dashed lines. The results show that controlling for the effects of ENSO and volcanic activity increases the estimate of the AF considerably, resulting in $\hat{\alpha}_{3}=47.16\%$ and $\hat{\alpha}_{4}=46.97\%$. The estimates of the standard deviations of the error terms $(\widehat{SD}(u_{t}^{(3)})=0.09$ and $\widehat{SD}(u_{t}^{(4)})=0.63)$ decrease substantially compared to those of models (1) and (2), indicating that the covariates ENSO and VAI explain much variation in the data. This is corroborated by the coefficient of determination ($R^{2}$) values reported in Table 1.444The $R^{2}$ of model (1) equals zero by construction, since this model only features an intercept. The decreased variance of the residuals from models (3) and (4) imply that their estimates of the constant AF $\alpha$ are more precise. The standard error of the estimate $\hat{\alpha}_{4}$ is approximately $16\%$ lower than the one of estimate $\hat{\alpha}_{3}$ and approximately $34\%$ lower than the one of the conventional estimate $\hat{\alpha}_{1}$. Our preferred estimate $\hat{\alpha}_{4}$ results in an estimated AF of $47.0\%$ (±$\pm 1.1\%$; $1\sigma$) with an associated $95\%$ confidence interval of $[44.9\%,49.0\%]$.
@@
The approximate constancy of the AF over the historical period 1959–2022, as documented in the literature and confirmed by the cointegration analysis in this study, can be taken as the result of a near exponential growth in emissions and an approximately linear response of the carbon sinks to atmospheric concentrations (Raupach et al., 2014). In scenarios describing the future, for example when emissions are declining, the AF is expected to depart from constancy and may vary over time (e.g., Jones et al., 2013; Pressburger et al., 2023). This motivates the specification of a time-varying AF $\alpha=\alpha_{t}$.
@@
where $\alpha_{t}$ is a time-varying coefficient representing the fraction of emitted CO2 that remains in the atmosphere in year $t$, i.e. the airborne fraction in year $t$. The ratio $G_{t}/E_{t}$ may be used to track the amount of emitted CO2 that remains airborne, and hence, the estimate $\hat{\alpha}_{1,t}=G_{t}/E_{t}$ can be regarded as appropriate, although very noisy. A possible solution to this problem is to apply a local smoothing operation, e.g. a two-sided moving average filter. In any case, filtered or not, the variability of the ratio $G_{t}/E_{t}$ will be amplified when future emissions $E_{t}$ start to approach zero.
@@
$G_{t}=\alpha_{t}E_{t}+u_{t}^{(2)}.$
A versatile way of treating such a time-varying regression model is to assume random walk dynamics for $\alpha_{t}$ and estimate it by means of a recursive regression filter, such as the Kalman filter and smoother (Durbin and Koopman, 2012, §3.6.1, §4.3, §4.4). This approach yields the minimum mean-squared error estimator $\hat{\alpha}_{2,t}$, and it does not suffer from the deficiencies of the time-varying ratio-based estimator $\hat{\alpha}_{1,t}$. Section S4 of the Supplementary Information provides further technical details on this procedure.
@@
As expected, the ratio-based estimate $\hat{\alpha}_{1,t}=G_{t}/E_{t}$ (blue) is a very noisy series, especially when $E_{t}\approx 0$. In contrast, the regression-based estimate $\hat{\alpha}_{2,t}$ (red) evolves over time in a stable fashion and shows sensible AF estimates, also when $E_{t}\approx 0$. A further benefit of the regression-based method is its delivery of confidence intervals for $\hat{\alpha}_{2,t}$ (shaded red area) which are not immediately available for the ratio-based estimator $\hat{\alpha}_{1,t}$. The covariates for El Niño and volcanic activity can readily be incorporated into the regression-based framework.
In the SSP1-2.6 scenario studied here, the regression-based AF estimate $\hat{\alpha}_{2,t}$ remains roughly constant until 2050, after which it gradually declines towards zero. In 2060 the atmospheric changes turn negative, resulting in a negative estimate of the AF, meaning that the sink uptake exceeds the emissions. Around 2070, the emissions turn negative as well, resulting in a positive estimate of the AF. After a short adjustment period, the estimate of the AF exceeds one, indicating that the sinks continue to absorb CO2 even in this regime with highly negative emissions.
Figure 2: Analysis of SSP1-2.6 data over the period 20232023202320232100210021002100. a) Atmospheric concentration changes ($G_{t}$) for the historical period 1959–2022 (black) and the SSP period 2023–2100 (blue, magenta). b) Emissions ($E_{t}$) data. Magenta lines show the output from MAGICC; blue lines show the perturbed data. c) Ratio of atmospheric changes to emissions ($G_{t}/E_{t}$). The red line in c) is the regression-based estimator $\hat{\alpha}_{2,t}$ of the time-varying airborne fraction $\alpha_{t}$, obtained from the Kalman smoother. Shaded area is a $95\%$ confidence band around $\hat{\alpha}_{2,t}$.

### v4

@@ Abstract
Previous studies have analyzed the AF as the ratio of yearly changes in atmospheric CO2 concentrations ($G_{t}$, numerator) and anthropogenic CO2 emissions ($E_{t}$, denominator) (e.g. Canadell et al., 2007; Raupach et al., 2008; Le Quéré et al., 2009; Knorr, 2009; Ballantyne et al., 2012; Raupach et al., 2014; Ballantyne et al., 2015; Keenan et al., 2016; Bennedsen et al., 2019; van Marle et al., 2022; Pressburger et al., 2023). An alternative to this approach is to consider the cumulative airborne fraction (CAF; e.g., Jones et al., 2013; Jones et al., 2016; Liddicoat et al., 2021), but since this approach is less commonly used in the literature and is less amenable to statistical analysis, we only briefly address the CAF in our empirical applications and omit it from our theoretical discussions. Instead, we follow the main body of the literature and adopt the conventional estimator of the AF which is defined as the sample mean of the yearly ratio $G_{t}/E_{t}$. In this paper, we show that this estimator suffers from a number of statistical deficiencies due to its definition as the ratio of two stochastic processes. One problem is that the denominator in the ratio, $E_{t}$, may have a positive probability density at zero, implying that the estimator does not possess any moments. This lack of moments implies, for instance, that the mean and variance of the estimator do not exist. Although this issue may be of limited concern during the period 1959–2022, it becomes important in future scenarios where CO2 emissions decrease, such as those consistent with “net-zero” CO2 emissions, a committed goal of the international community (Rockström et al., 2017; Riahi et al., 2022). For example, issues with the conventional analysis of the AF when emissions decrease are recently encountered in Pressburger et al. (2023), where the future AF implied by output from a climate model is studied. Another problem is that the time series of yearly changes in atmospheric concentrations $G_{t}$ and emissions $E_{t}$ exhibit a trend, which implies that a central limit theorem does not hold for the estimator. Therefore, its limiting distribution may be non-Gaussian, unless a separate assumption of Gaussianity is imposed. This implies that confidence intervals and $p$-values for test statistics based on the Gaussian distribution may not be valid for the conventional ratio-based estimator.
@@
We show that the time series of yearly changes in atmospheric CO2 concentrations, $G_{t}$, cointegrates with the time series of anthropogenic CO2 emissions, $E_{t}$, over the period 1959–2022. Cointegration means that, even though $G_{t}$ and $E_{t}$ are non-stationary time series individually, the linear regression equation $G_{t}=\alpha E_{t}+u_{t}$, with constant parameter $\alpha$, has a stationary error $u_{t}$. In other words, the non-stationary processes $G_{t}$ and $E_{t}$ are cointegrated when a linear combination exists that yields a stationary process. The parameter $\alpha$ denotes the fraction of emissions that is added to the atmosphere in year $t$, i.e., in the linear regression, $\alpha$ represents the AF. We show the constancy of $\alpha$ statistically on the historical sample. In future scenarios where emissions decrease substantially, $\alpha$ will be time-varying, however.
@@
Figures 1a) and 1b) show yearly changes in atmospheric CO2 ($G_{t}$) and yearly CO2 emissions from anthropogenic sources ($E_{t}$), respectively. The black line in Figure 1c) shows the ratio of these two variables, $G_{t}/E_{t}$. Data are obtained from the Global Carbon Project and cover the period 19591959195919592022202220222022; in the empirical section below we provide further details on the data. The most conspicuous feature of the two data series $G_{t}$ and $E_{t}$ is that both time series exhibit upwards trends. Trending behavior is indicative of time series being non-stationary. A simple least-squares statistical analysis of the bivariate system $(G_{t},E_{t})$, where the non-stationarity of a time series is not accounted for, yields invalid inference and should be avoided (Granger and Newbold, 1974). However, the notion of cointegration (see, for example, chapter 19 in Hamilton, 1994, for a textbook treatment) allows us to keep working with the trending time series $G_{t}$ and $E_{t}$ while still obtaining valid statistical inference. Cointegration methods have been applied in earlier climate studies (e.g. Kaufmann and Stern, 2002; Schmith et al., 2012). Informally, two time series are cointegrated if they share a common trend. Formally, the time series $G_{t}$ and $E_{t}$ are said to be cointegrated when both $G_{t}$ and $E_{t}$ are non-stationary, and the error term $u_{t}$ in the regression equation $G_{t}=\alpha E_{t}+u_{t}$ is stationary. We adopt the Dickey-Fuller test (Dickey and Fuller, 1979) to determine whether a time series is non-stationary. The test statistic is for the null hypothesis of a unit root, that is, of having a unit value for the autoregressive dependence of $G_{t}$ on its lagged value $G_{t-1}$ (and $E_{t}$ on $E_{t-1}$). The results from this test strongly suggest that both time series are non-stationary (Supplementary Information S2.1). The null hypothesis of no cointegration (that is, $u_{t}$ is non-stationary) can be tested formally using the Engle-Granger test (Engle and Granger, 1987), which is a Dickey-Fuller test on the residuals in the regression $G_{t}=\alpha E_{t}+u_{t}$, adjusted for the fact that these residuals are not observed but must be estimated. The null hypothesis of a unit root in $u_{t}$ is firmly rejected (Supplementary Information S2.1), and we can therefore conclude that the two yearly time series $G_{t}$ and $E_{t}$ are cointegrated.
@@
for $t=1,2,\ldots,T$, where the disturbance $u^{(2)}_{t}$ is also a zero-mean error process. A model closely related to (2) has previously been used to reconstruct and predict CO2 growth rates (Jones and Cox, 2005; Betts et al., 2016). The relationship between the disturbances in equations (1) and (2) is given by $u^{(1)}_{t}=u^{(2)}_{t}\,/\,E_{t}$. Cointegration of $G_{t}$ and $E_{t}$ implies that $u_{t}^{(2)}$ is a stationary process. Then, the parameter $\alpha$ can be estimated directly using a simple least-squares calculation, yielding the regression-based estimator as given by
@@
Table 1: Least-squares regression output of the linear models (1)–(4) for the Global Carbon Project data on the full sample 1959–2022 (left panel) and the subsample 1992–2022 (right panel). Estimates of the AF $\alpha$ are denoted by “$\hat{\alpha}$”. The table reports standard errors “SE($\hat{\alpha}$)” and standard errors relative to SE($\hat{\alpha}_{1}$) from model (1). Relative SE less than one indicates lower estimation uncertainty than the ratio-based estimator. The $95\%$ confidence interval for the AF $\alpha$, based on the Gaussian distribution, is denoted by “$CI_{95\%}(\alpha)$”, the estimated standard deviation of the error $u_{t}$ by “$\widehat{SD}(u_{t})$”, and the coefficient of determination by “$R^{2}$”.
@@
By introducing covariates in models (1) and (2), we can reduce the variance of the error processes $u_{t}$ and thus achieve more precise estimates of the AF $\alpha$. For example, it is common practice in the literature to control for the effects of El Niño ($ENSO_{t}$) and volcanic activity ($VAI_{t}$), (e.g. Raupach et al., 2008; van Marle et al., 2022). [^3] VAI data for volcanic activity are obtained from Ammann et al. (2003). ENSO data are constructed from the Niño 3 SST Index of the National Oceanic and Atmospheric Administration (NOAA), available at https://psl.noaa.gov/gcos_wgsp/Timeseries/Data/nino3.long.anom.data, last accessed June 17, 2024. We have converted monthly ENSO data into a yearly time series of September-August ENSO means (Jones et al., 2001). This 4-month lag provides the best fit between $G_{t}$ and $ENSO_{t}$. The slight trend in yearly ENSO data is removed so that it has no impact on the AF estimates. The data are shown in Figure S2 of the Supplementary Information. Hence, we consider the models
@@
for $t=1,2,\ldots,T$, where $\tilde{\gamma}_{i}$ and $\gamma_{i}$, for $i=1,2$, are regression coefficients, and the model errors $u^{(j)}_{t}$ follow a zero-mean error process, for $j=3,4$. For both models, the coefficients can be estimated using least-squares regression. Let $\hat{\alpha}_{3}$ and $\hat{\alpha}_{4}$ denote the least-squares estimators of $\alpha$ from models (3) and (4), respectively. The estimation results relevant for the AF $\alpha$ are presented in columns labeled as (3) and (4) of Table 1, while the estimation results for the ENSO and VAI coefficients $\gamma_{i}$ and $\tilde{\gamma}_{i}$, for $i=1,2$, are reported in Table S2 of the Supplementary Information. The regression fits of models (3) and (4) are presented as in Figure 1c) and 1e) and their associated estimated residuals $\hat{u}_{t}$ in Figure 1d) and 1f), all as red dashed lines. The results show that controlling for the effects of ENSO and volcanic activity increases the estimate of the AF considerably, resulting in $\hat{\alpha}_{3}=47.16\%$ and $\hat{\alpha}_{4}=46.97\%$. The estimates of the standard deviations of the error terms $(\widehat{SD}(u_{t}^{(3)})=0.09$ and $\widehat{SD}(u_{t}^{(4)})=0.63)$ decrease substantially compared to those of models (1) and (2), indicating that the covariates ENSO and VAI explain much variation in the data. This is corroborated by the coefficient of determination ($R^{2}$) values reported in Table 1.444The $R^{2}$ of model (1) equals zero by construction, since this model only features an intercept. The decreased variance of the residuals from models (3) and (4) imply that their estimates of the constant AF $\alpha$ are more precise. The standard error of the estimate $\hat{\alpha}_{4}$ is approximately $16\%$ lower than the one of estimate $\hat{\alpha}_{3}$ and approximately $34\%$ lower than the one of the conventional estimate $\hat{\alpha}_{1}$. Our preferred estimate $\hat{\alpha}_{4}$ results in an estimated AF of $47.0\%$ (±$\pm 1.1\%$; $1\sigma$) with an associated $95\%$ confidence interval of $[44.9\%,49.0\%]$. The slightly increased AF estimates for models (3) and (4), compared to the models (1) and (2) without covariates, confirm a similar finding in Betts et al. (2016).
@@
The approximate constancy of the AF over the historical period 1959–2022, as documented in the literature and confirmed by the cointegration analysis in this study, can be taken as the result of a near exponential growth in emissions and an approximately linear response of the carbon sinks to atmospheric concentrations (Raupach et al., 2014). In scenarios describing the future, for example when emissions are declining, the AF is expected to depart from constancy and may vary over time (e.g., Jones et al., 2013; Pressburger et al., 2023). This motivates the specification of a time-varying AF $\alpha=\alpha_{t}$, with $\alpha_{t}$ denoting the AF in year $t$, i.e. the fraction of emissions ($E_{t}$) added to the atmosphere ($G_{t}$) in year $t$.
@@
where $\alpha_{t}$ is a yearly time-varying coefficient. The ratio $G_{t}/E_{t}$ may be used to track the amount of emitted CO2 that remains airborne, and hence, the estimate $\hat{\alpha}_{1,t}=G_{t}/E_{t}$ can be regarded as an appropriate but very noisy time-varying AF. A possible way to reduce the noise in AF is to apply a local smoothing operation, e.g. a two-sided moving average filter. In any case, filtered or not, the variability of the ratio $G_{t}/E_{t}$ will be amplified when future emissions $E_{t}$ start to approach zero. Another possible solution for noise reduction previously suggested in the literature, is to use the cumulative AF (CAF) in place of the yearly AF (e.g., Jones et al., 2013; Jones et al., 2016; Liddicoat et al., 2021). However, due to its cumulative nature, the CAF can be slow to detect changes in the behavior of the carbon sinks, making it less useful for the purpose of analyzing a time-varying AF (Supplementary Information S3).
@@
$G_{t}=\alpha_{t}E_{t}+u_{t}^{(2)}.$ (5)
A versatile way of treating such a time-varying regression model is to assume random walk dynamics for $\alpha_{t}$ and estimate it by means of a recursive regression filter, such as the Kalman filter and smoother (Durbin and Koopman, 2012, §3.6.1, §4.3, §4.4). This approach yields the minimum mean-squared error estimator $\hat{\alpha}_{2,t}$, and it does not suffer from the deficiencies of the time-varying ratio-based estimator $\hat{\alpha}_{1,t}$. In the regression-based model (5), $\alpha_{t}$ is multiplied by $E_{t}$. When emissions turn negative for the first time in some year $t$, i.e. $E_{t}<0$, we let $\alpha_{t}$ be reflected around one. For this purpose, we adjust the random walk specification at year $t$ with a one-time instantaneous shift in $\alpha_{t}$ when $E_{t}<0$ for the first time. Section S5 of the Supplementary Information provides further motivation and technical detail on this procedure.
@@
The ratio-based estimate $\hat{\alpha}_{1,t}=G_{t}/E_{t}$ (blue) is a very noisy series as we have anticipated, especially when $E_{t}\approx 0$. In contrast, the regression-based estimate $\hat{\alpha}_{2,t}$ (red) evolves over time in a stable fashion and shows sensible AF estimates, also when $E_{t}\approx 0$. A further benefit of the regression-based method is its delivery of confidence intervals for $\hat{\alpha}_{2,t}$ (shaded red area) which are not immediately available for the ratio-based estimator $\hat{\alpha}_{1,t}$. The covariates for El Niño and volcanic activity can readily be incorporated into the regression-based framework.
In the SSP1-2.6 scenario studied here, the regression-based AF estimate $\hat{\alpha}_{2,t}$ remains roughly constant until 2050, after which it gradually declines towards zero. In 2060, the atmospheric changes turn negative, resulting in a negative estimate of the AF, meaning that the sink uptake exceeds the emissions. In 2077, the emissions turn negative as well, instigating the visible switch to a positive estimate of the AF. The estimates of the AF exceed one from 2077 onwards, indicating that the sinks continue to absorb CO2 even in this regime with highly negative emissions. These findings can be contrasted with the analysis of the SSP1-1.9 scenario (Supplementary Information, Figure S5) which has a similar trajectory for the regression-based AF estimates $\hat{\alpha}_{2,t}$ as for the SSP1-2.6 scenario, except that from 2080 onwards we have $\hat{\alpha}_{2,t}<1$ implying that the sinks turn into carbon sources (releasing more carbon dioxide than they absorb).
Figure 2: Analysis of SSP1-2.6 data over the period 20232023202320232100210021002100. a) Atmospheric concentration changes ($G_{t}$) for the historical period 1959–2022 (black) and the SSP period 2023–2100 (blue, magenta). b) Emissions ($E_{t}$) data. Magenta lines show the output from MAGICC; blue lines show the perturbed data. c) Ratio of atmospheric changes to emissions ($G_{t}/E_{t}$). The red line in c) is the estimated fraction of emissions ($E_{t}$) added to the atmosphere ($G_{t}$) in each year $t$, and, more specifically, it is the regression-based estimator $\hat{\alpha}_{2,t}$ of the time-varying airborne fraction $\alpha_{t}$, obtained from the Kalman smoother. Shaded area is a $95\%$ confidence band around $\hat{\alpha}_{2,t}$.